OpenStudy (anonymous):

what is the value of x? given: 6x+y=9.7 4x+2y=9.8 Please show me the steps...Thanks

OpenStudy (anonymous):

8x=9.7*2-9.8=9.6 x=1.2

OpenStudy (smurfy14):

y=2.5 x=1.2

OpenStudy (smurfy14):

first prob is y=-6x+9.7 then put tht in ur second equation 4x+2(-6x+9.7)=9.8 then distribute and solve from x from there. after you get x=1.2 go back and plug 1.2 from x in original equation and solve

OpenStudy (anonymous):

we have given 2 equations. 6x + y = 9.7..............................(1) 4x + 2y = 9.8..........................................(2) first eliminate y, by multiplying equation (1 )by 2 12x + 2y = 19.4................................(i) 4x + 2y = 9.8.............................(ii) ............................................................................on subtracting (i) and (ii) we get 8x = 9.6 divide both side by 8 we have x = 1.2 plug the value of x in 1 we get 6(1.2) +y =9.7 y = 9.7 - 7.2 y= 2.5 So , anwser is x=1.2 and y = 2.5